An electron starts from rest 37.0cm from a fixed point charge with Q = -0.105?C .

How fast will the electron be moving when it is very far away?
Answer #1

Ei = ef

k e Q/r = 1/2 mv^2

9.0E9*1.6E-19*0.105E-6/.37 = 0.5*9.11E-31*v^2

v=2.99E7 m/s
Answer #2

Ei = ef

k e Q/r = 1/2 mv^2

9.0E9*1.6E-19*0.105E-6/.27 = 0.5*9.11E-31*v^2

v=3.51E7 m/s